OpenStudy (anonymous):

Solve the problem Helpp plzzzz An initial investment of $14,000 is appreciated for 12 years in an account that earn 10% interest, compounded semiannually . find the amount of money in the account at the end of the period

OpenStudy (mathstudent55):

\(F = P(1 + \dfrac{r}{n})^{nt} \) F = future value P = present value r = annual interest rate written as decimal n = number of compounding periods in one year t = number of years

OpenStudy (anonymous):

I dont got it

OpenStudy (cwrw238):

well plug in the values in question to the formula that math student gave you eg: P = 14,000 r = 0.10 ( 10% written as a decimal) can you tell me what n and t are?

OpenStudy (anonymous):

n= number of compounding periods in one year t = number of years

OpenStudy (cwrw238):

ok and what is the values of these ( check the question)

OpenStudy (anonymous):

n=? t=12

OpenStudy (cwrw238):

n = 2 ( semiannually means 2 times a year)

OpenStudy (cwrw238):

so plug these values in and use ur calculator to find the value of F

OpenStudy (cwrw238):

s its F = 14000(1 + ( 0.10/2))^24

OpenStudy (anonymous):

how did u get 24

OpenStudy (anonymous):

i'm confused

OpenStudy (cwrw238):

nt - that is n times t = 2*12 = 24

OpenStudy (anonymous):

i got it now

OpenStudy (cwrw238):

good

OpenStudy (anonymous):

the answer will be 14000

OpenStudy (cwrw238):

no you work out 14000(1 + 0.05)^24 using your calculator

OpenStudy (cwrw238):

the amount at end of the period = F which is given by the formula

OpenStudy (anonymous):

45151.4

OpenStudy (anonymous):

right

OpenStudy (mathstudent55):

\(F = P(1 + \dfrac{r}{n})^{nt}\) \(P = $14,000\) \(r = 10\% = 0.1\) \(n = 2\) \(t = 12\) \(F = $14,000 \left(1 + \dfrac{0.1}{2} \right)^{2 \times 12} \) \(F = $45,151.40\)